OpenStudy (anonymous):

∜625 I DONT KNOW HOW TO DO THIS equation

OpenStudy (anonymous):

ive been trying to do this for the last for the past 30 mins ._.

OpenStudy (ipwnbunnies):

Can you factor 625?

OpenStudy (ipwnbunnies):

Btw, we're not solving it, I guess. We're simplifying it.

OpenStudy (anonymous):

\[625=5^4\] so \[\sqrt[4]{625}=5\]

OpenStudy (anonymous):

OMG thank you your so much help :)

OpenStudy (mrnood):

since you now have th eanswer - are you interested in a method?

OpenStudy (anonymous):

Yes please!!!

OpenStudy (anonymous):

because the next one is the same thing except with an x-variable in the problem ._.

OpenStudy (mrnood):

OK \[x= \sqrt[4]{625}=625^{\frac{ 1 }{ 4}}\] so take logs \[logx= \log\sqrt[4]{625}=\log625^{\frac{ 1 }{ 4}}\] \[logx= \frac{ 1 }{ 4 }\log625\] So oyu can solve x fro any root using logs

OpenStudy (anonymous):

Thank you so much that is a great help i will be sure to write it down in my notebook for when i need it next time :)